# RDA-PSO: A computational method to quantify the diffusive dispersal of insects

### v1

@@ 4.4 Comparison of TC, ATC, MDT, and RDA-PSO estimates
Hainan - center 12.81 36.89 32.10 40.78
Hainan - edge 109.36 119.4 201.87 155.48
Cairns 190.31 256.65 243.79 436.84
@@
Whereas $D_{RDA}$ and $D_{ATC}$, and separately $D_{RDA}$ and $D_{MDT}$, are reasonably close to one another for Hainan, there is a big difference in the two estimates for Cairns. In the latter case, as well as for the Hainan center experiment, $D_{MDT}$ and $D_{ATC}$ provide similar estimates. In addition, $D_{ATC}$ always under-estimates $D_{RDA}$. In all cases, $D_{TC}$ is much smaller than the other estimates. The latter illustrates how not compensating for the lack of spatial uniformity in the distribution of the collection sites affects estimates of mosquito dispersal.
@@ 5 Discussion
This article introduces the RDA-PSO method, which is a novel computational approach that efficiently estimates the diffusion coefficient $D$ associated with mark-release-recapture (MRR) experiments. The only assumption is that the dispersal of insects in those experiments is diffusive. Input data consist of temporal and spatial recapture ratios (defined in Section 3.2) and the output is an estimate of $D$, called $D_{RDA}$. The method itself (see Section 3.6 as well as Appendices B and C) involves a forward problem in which independent walkers diffuse from a release point and may be recaptured at collection sites whose locations match those used in the MRR experiment. Spatial and temporal recapture ratios obtained for each set of random walk (k) and trap (q, p, and ${\tt s}_{\tt e}$) parameters are compared to the observed MRR ratios to define an error, which is minimized in the parameter space by a particle swarm optimization (PSO) method (Section 3.6.2 and Appendix C). The result is an estimate of the values of k, q, p, and ${\tt s}_{\tt e}$ that best match the observed spatial and temporal recapture ratios. The estimate of k is then converted into an estimate of the diffusion coefficient $D$.
@@
We also introduced two other optimization approaches to estimate $D$, based on theoretical considerations: the time-corrected model (Section 3.3), valid when recapture sites are uniformly distributed in the study area, and the area-and-time-corrected model (Section 3.4), useful for unevenly distributed collection sites. We explored how these estimates vary with trap location, trap density, and frequency of insect collection. More importantly, we used the time-corrected and area-and-time-corrected models as baselines against which we compared the new computational method introduced here. A third way to estimate $D$, based on the empirical MDT (Section 3.5), provides results that are sometimes larger and other times smaller than those of the area-and-time-corrected model.
@@
Through systematic testing, we observed that, when applied to synthetic data, $D_{RDA}$ is a very accurate estimate of the actual diffusion coefficient used to generate the data. We believe this is due to the existence of a unique minimizer in the explored parameter area, whose location is first estimated through a grid search and then refined by the PSO method (see Appendix D for an example in the $({\tt k},{\tt q})$ plane). Similar performance was not observed for $D_{TC}$, which required an increase in collection frequency to become accurate, even for uniformly distributed traps (Appendix G.1). In addition, we found that, on average $D_{ATC}$ tends to overestimate $D$ due to round-off errors associated with the application of the correction factor (Appendix G.2).
When applied to real-world data (Section 4.3 and Appendix E), which typically involves unevenly distributed collection locations, $D_{ATC}$ tends to be lower than $D_{RDA}$. This suggests that estimates of $D$ based on theoretical calculations that rely on a Gaussian probability distribution function tend to undervalue $D$, and consequently undervalue estimates of an insect MDT over a given period of time. Similarly, estimates of $D$ based on the empirical MDT also seem to be unreliable since they either underestimate or overestimate insect dispersal properties. For mosquitoes like Aedes aegypti, which may transmit life-altering diseases such as dengue, Zika, and chikungunya, underestimating how far they may disperse once infected may have serious public health consequences.
In contrast, the computational method presented here produces estimates of $D$ that are robust to changes in the density and location of collection sites, and does not require adjusting the observed recapture ratios, even if collection sites are unevenly distributed in the study area. Moreover, reliable estimates of k may easily be obtained via repeated application of the RDA-PSO method (as seen in Figure 1) and appear to be normally distributed in the parameter space (Figure 2). This in turn makes it relatively easy to approximate a confidence interval for $D_{RDA}$, based on different applications of the computational method.
@@
We therefore believe the computational method presented here provides a novel and effective way to analyze MRR data and subsequently determine insects dispersal properties in the corresponding environment. As reviewed in Section 2, a diffusive approximation to describe the spread of insects is appropriate and commonly used in partial differential equation (PDE) compartmental models describing the growth and spread of mosquitoes. The proposed approach improves on estimates of $D$ based on theoretical considerations, which often require frequent sampling and involve correcting recapture data according to how collection sites are distributed in the study area. As previously mentioned, such adjustments are not necessary for the application of the RDA-PSO method, whose estimates of $D$ are not constrained by the density, nature, and location of collection sites, or by trap collection frequency. Consequently, the present method may be used not only to inform which range of values of $D$ should be included in PDE models, but also to estimate the lifetime MDT of mosquitoes like Aedes aegypti in different environments, together with the associated risk of disease spread by these vectors. A complete analysis of the MRR literature and the estimation of associated dispersal parameters based on the RDA-PSO method introduced here is left for future work.
@@ Acknowledgements
This material is partly based upon High Performance Computing (HPC) resources supported by the University of Arizona TRIF, UITS, and Research, Innovation, and Impact (RII) and maintained by the UArizona Research Technologies department. Research reported in this publication was supported in part by the University of Arizona’s BIO5 Institute Team Scholars Program. Mount Holyoke College undergraduate students, Chuhan Wang ‘24 and Karry Wang ‘23, worked on an image processing project in Summer 2022, which helped with the digitization of maps used in this work. The authors would like to thank Heidi Brown for helpful conversations on mosquito behavior.
@@ Appendix A Diffusion Equation
As described in Section 2, diffusive mosquito dispersal means that at the microscopic level, the two-dimensional (ground) projection of the flight path of a moving mosquito consists of an alternation of straight flight segments of fixed length and angular reorientations uniformly distributed in $[0,2\pi]$. At the macroscopic level, this corresponds to planar diffusion, described by the partial differential equation
@@ B.2 Processing: Release and recapture of diffusive agents
$=\zeta\left(1.01\max_{\Omega_{c}}||\nabla F||+S_{x}\cos\phi+S_{y}\sin\phi\right),$ (17)
where $\zeta$ is a positive constant that will be determined below, $\phi$ is the angle of motion of the walker measured from the $x$-axis, and $\Omega_{c}$ represents all of the locations inside any capture site in the simulation. The quantity ${\tt dBase}=1.01\max_{\Omega_{c}}||\nabla F||$ is chosen to be slightly bigger than $\max_{\Omega_{c}}||\nabla F||$, so that the walker always changes position, that is, d $>0$. A simple calculation shows that $||\nabla F||$ is maximal on a circle of radius $r=1/\sqrt{2\alpha}$ centered on each capture site, and that
$\max_{\Omega_{c}}||\nabla F(x,y)||=\frac{\sqrt{2}}{\pi}\alpha^{3/2}e^{-1/2}.$
The expression for d is smaller when the walker is heading away from the center of the capture site, and larger if it is heading towards it. A plot of d$/\zeta$ as a function of $\phi$ for a walker located at $(2,0)$ is shown in the left panel of Figure 3. Since the center of the capture site is at $(0,0)$, an angle $\phi=\pi$ indicates the walker moves towards the center of the capture site while $\phi=0$ or $\phi=2\pi$ indicates it is moving away from the capture region.
Figure 3: Left panel: Plot of d/$\zeta$ (see Equation (17)) as a function of $\phi$ for a walker located at $(x,y)=(2,0)$, assuming the trap center is at $(x_{0},y_{0})=(0,0)$. We let $\alpha=0.2$, which leads to dBase = 0.0247. Top right panel: Range of the step sizes d as a function of the distance $r$ between the walker location and the center of a capture site. The vertical segment at OR/3 shows the range with largest extent. The exact value of d depends on the angle $\phi$. Bottom right panel: Possible steps taken within a trap of outer radius OR = 10. The red dots represent a walker’s current location. The blue circles around each red dot visualize how far the walker would move if it were to head in that specific direction
The length of a step taken by a walker within a capture site needs to be small compared to the size of the site, so that the walker does not immediately leave the site upon entering it. To this end, we choose $\zeta$ so that $\zeta(2.01\max_{\Omega_{c}}||\nabla F(x,y)||)={\tt OR}/3$, which guarantees that ${\tt d}\leq{\tt OR}/3$. This gives, after substituting $\alpha$ by its chosen value in terms of OR from above,
@@ Appendix C Particle Swarm Optimization
and $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are 4-dimensional vectors of independent random numbers drawn from the uniform distribution in the interval $[0,1]^{4}$. Note that “*” here denotes element-wise multiplication. The parameters are $w$, the “inertia” weight, $\phi_{p}$, the “cognitive” coefficient, and $\phi_{g}$, the “social” coefficient. These heuristic coefficients are chosen so that $10<w<$ 1 and $\phi_{p},\phi_{g}\geq 1$, reflecting a balance between moving towards the “best seen” position and swarming for a more effective search. In particular, with $w<1$, the first term acts to slow down the particle when getting close to a minimum. The remaining two terms move it generally in the direction of its own local best position (second term) and the swarm’s global best position (third term), with an element of randomness included in that motion. We initiate $\mathbf{v}$ randomly in $[-1,1]^{4}$, and denote the resulting vector from the previous step by $\mathbf{v}_{old}$. For all of the simulations presented here, we set $w=0.4$ and $\phi_{p}=\phi_{g}=1.2$. A new generation of particles is obtained after all of the positions of the previous generation have been updated once according to the above formula.
@@ Appendix D Virtual MRR experiment
This section illustrates how the different methods introduced in this article perform on a virtual MRR experiment. To this end, we first obtain synthetic capture ratios by simulating the forward problem with fixed parameters and averaging the resulting ratios of 100 simulations.
@@ D.2 Generation of MRR data for the virtual experiment
In order to limit the effect of randomness in the generated data, we run 100 simulations of the forward problem with the above setup and use average recapture ratios as the observed ratios. In each of the 100 simulations, $N=$100,000 agents are released from the middle of the grid. Each moves independently of the others and performs a random walk with parameters k = 100 and q = 30 for a total duration of 20 days. Trap efficiency is set at 100% and p = 0. Table 4 lists the average capture ratios, by day (with a count every other day) and by zone, resulting from these simulations. The selected value of k corresponds to a diffusion coefficient $D=300$ m2 per day.
@@
D.3 Application of the time- and area-and-time-corrected models
D.4 Application of the MDT model
For completion, we calculate the empirical MDT by multiplying the spatial ratios $\sigma_{j}$ element-wise by the average distance to each corresponding zone, that is

### v2

@@ 4.4 Comparison of MDT, TC, ATC, and RDA-PSO estimates
Hainan - center 32.10 12.81 36.89 40.78
Hainan - edge 201.87 109.36 119.4 155.48
Cairns 243.79 190.31 256.65 436.84
@@
Fitting the temporal and spatial capture ratios reported in each of Tables 3 and 4 into the time-corrected model (10) leads to estimates of the diffusion coefficient shown in the second column ($D_{TC}$) of Table 5. Since capture sites are not uniformly distributed across collection zones, we also show in column 3 ($D_{ATC}$) of Table 5 the area-and-time-corrected estimates of $D$. We note however that, contrary to the Cairns study, there is not sufficient information in the Hainan data [80] to correct the temporal ratios after having applied the correction factor to the spatial ratios. In other words, recapture numbers are reported per zone and per day in [63] but not in [80]. For comparison, the first column lists estimates of $D$ based on averages of reported values of the MDT: between 10 and 20 m for females released at the center of the village and close to 40 m for those released at the periphery in Hainan [80]; between 77.6 and 77.8 m in Cairns [63]. The last column shows estimates of $D$ obtained by applying the RDA-PSO method.
Whereas $D_{RDA}$ and $D_{ATC}$, and separately $D_{RDA}$ and $D_{MDT}$, are reasonably close to one another (with a relative difference less than 30%) for Hainan, there is a big difference in the two estimates for Cairns. In the latter case, as well as for the Hainan center experiment, $D_{MDT}$ and $D_{ATC}$ provide similar estimates. In addition, $D_{ATC}$ always under-estimates $D_{RDA}$. In all cases, $D_{TC}$ is much smaller than the other estimates. The latter illustrates how not compensating for the lack of spatial uniformity in the distribution of the capture sites affects estimates of mosquito dispersal.
A simple intuitive interpretation of the differences in the estimates of the diffusion coefficient can be obtained using the first equation in (3), which approximates the average diffusive displacement per day. For example, in the Hainan center experiment (see Table 5), the time-corrected estimate implies an average displacement of 6.34 m/day, whereas the RDA estimate results in a displacement of 11.32 m/day. For the Cairns experiment, the MDT estimate corresponds to a displacement of 27.67 m/day, while the RDA estimate amounts to a displacement of 37.05 m/day. Although most of these estimates are of the same order of magnitude, and are consistent with the typical flight range of Ae. aegypti ($\lesssim 30$ m/day [36]), there are notable differences in their specific values.
@@ 5 Discussion
This article introduces the RDA-PSO method, which is a novel computational approach that efficiently estimates the diffusion coefficient $D$ associated with mark-release-recapture (MRR) experiments. The only assumptions we make are that the dispersal of insects in those experiments is diffusive (with bias in the attracting region of the capture sites), and that MRR experiments are conducted in the absence of advection, in an environment where most of the geographic heterogeneity is associated with the presence of capture sites (houses or traps). Input data consist of temporal and spatial recapture ratios (defined in Section 3.1.2) and the output is an estimate of $D$, called $D_{RDA}$. The method itself (see Section 3.5 as well as Appendices B and C) involves a forward problem in which independent walkers diffuse from a release point and may be recaptured at sites whose locations match those used in the MRR experiment. Spatial and temporal recapture ratios obtained for each set of random walk (k) and trap (q, p, and ${\tt s}_{\tt e}$) parameters are compared to the observed MRR ratios to define an error, which is minimized in the parameter space by a particle swarm optimization (PSO) method (Section 3.5.2 and Appendix C). The result is an estimate of the values of k, q, p, and ${\tt s}_{\tt e}$ that best match the observed spatial and temporal recapture ratios. Our numerical explorations show that although the parameters q, p, and ${\tt s}_{\tt e}$ are not individually identifiable, the parameter k is. The latter is then converted into an estimate of the diffusion coefficient $D$.
@@
We also introduced two other optimization approaches to estimate $D$, based on theoretical considerations: the time-corrected model (Section 3.3), valid when recapture sites are uniformly distributed in the study area, and the area-and-time-corrected model (Section 3.4), useful for unevenly distributed capture sites. We explored how these estimates vary with trap location, trap density, and frequency of insect collection. More importantly, we used the time-corrected and area-and-time-corrected models as baselines against which we compared the new computational method introduced here. A third way to estimate $D$, based on the empirical MDT (Section 3.2), provides results that are sometimes larger and other times smaller than those of the area-and-time-corrected model.For mosquitoes like Aedes aegypti, which may transmit life-altering diseases such as dengue, Zika, and chikungunya, underestimating how far they may disperse once infected may have serious public health consequences.
The four methods introduced here rely on spatial and temporal ratios (defined in Equations (4) and (5)). The temporal ratios $\tau_{i}$ may be viewed as weights that correct the formulas obtained from the continuous diffusion assumption in the MDT, TC, and ATC models (see Equations (6) and (10)). By definition, these weights are proportional to the daily number of recaptured mosquitoes, which reflects pool reduction due to previous recapture, death, or dispersal beyond the area of interest. Matching the $\tau_{i}$’s during the optimization process provides a simple way to factor in such depletion, without having to make assumptions regarding its specific causes.
@@
Through systematic testing, we observed that, when applied to synthetic data, $D_{RDA}$ is a very accurate estimate of the actual diffusion coefficient used to generate the data. We believe this is due to the existence of a unique minimizer in the explored parameter area, whose location is first estimated through a grid search and then refined by the PSO method (see Appendix D for an example in the $({\tt k},{\tt q})$ plane). Similar performance was not observed for $D_{TC}$, which required an increase in collection frequency to become accurate, even for uniformly distributed traps (Appendix G.1 and Table 2). In addition, we found that, on average $D_{ATC}$ tends to underestimate $D$, although only correcting spatial ratios led on average to overestimates, due to round-off errors associated with the application of the correction factor (Appendix G.2). Similar results were observed with real-world data (Section 4.4), which typically involves unevenly distributed capture sites: $D_{ATC}$ tends to be lower than $D_{RDA}$. Separately, $D_{MDT}$ tends to be lower than $D_{ATC}$ (Appendix G.2 and Section 4.4), except for the Hainan - edge study (Table 5). In general, correcting the temporal and spatial ratios to account for uneven trap density produces values of $D_{MDT}$ and $D_{ATC}$ that are lower than partially corrected estimates that only include corrections on spatial ratios. Correcting temporal ratios thus improves $D_{ATC}$ estimates (they get closer to the known value of $D$) but worsens $D_{MDT}$ estimates. Depending on experimental design, the data collected in MRR experiments may not provide sufficient information to correct both temporal and spatial ratios. This is for instance the case for the Hainan study.
Such a difference in performance highlights an advantage of the computational approach: the RDA-PSO method works regardless of the sampling frequency (since it calculates the ratios accordingly in the forward model), it produces estimates of $D$ that are robust to changes in the density and location of capture sites, and it does not require adjusting the observed recapture ratios, even if capture sites are unevenly distributed in the study area. In other words, the intrinsic discrete nature of RDA-PSO presents a distinct advantage over theoretical methods such as the MDT, TC, and ATC models, which are based on continuous sampling assumptions and are thus limited by collection frequency and density of capture sites.
The RDA-PSO method is stochastic in nature, leading to variability in the optimal parameter values that it returns. This is compounded by the use of a relatively small number of swarming particles and generations to reduce computational cost. However, reliable estimates of k may easily be obtained via repeated application of the RDA-PSO method (as seen in Figure 1) and appear to be normally distributed in the parameter space (Figure 2). This in turn makes it relatively easy to approximate a confidence interval for $D_{RDA}$, based on different applications of the computational method. We therefore recommend running multiple (e.g. 10 or 20) simulations and removing outliers based on the returned error estimate, as discussed in Section 4.3.2. Even though variability may be observed in the returned values of q, p, and ${\tt s}_{\tt e}$, we reiterate that associated values of k, and thus the estimates of $D_{RDA}$, are not noticeably affected by the trap parameters, including the trap efficiency.
The computational cost of the method is discussed in Appendix C.2. Although the results of Figure 1 required High Performance Computing (HPC) to demonstrate that the estimates of $D_{RDA}$ were normally distributed, a reliable value of the diffusion coefficient may be obtained in less than an hour by running a dozen simulations on a laptop.
@@
We therefore believe the RDA-PSO method presented here provides a novel and effective way to analyze MRR data and subsequently determine insects dispersal properties in the corresponding environment. As reviewed in Section 2, a diffusive approximation to describe the spread of insects is appropriate and commonly used in partial differential equation (PDE) compartmental models describing the growth and spread of mosquitoes. The present method may be used not only to inform which range of values of $D$ should be included in PDE models, but also to estimate the lifetime MDT of mosquitoes like Aedes aegypti in different environments, together with the associated risk of disease spread by these vectors. A systematic analysis of the MRR literature and estimation of associated dispersal parameters based on the RDA-PSO method introduced here is left for future work.
The above discussion situates the RDA-PSO method as a useful tool that can be incorporated into a larger array of predictive models for vector control strategies. Our current work can indeed be extended in several directions. It is straightforward to adapt our method to quantify other types of biological dispersal, as long as the process is diffusive and MRR or similar data are available. Furthermore, our code can be modified to incorporate local features such as constant directional biases (due for instance to the presence of wind or corridors such as streets in a city), attracting spots that are not capture sites and have different levels of attractiveness, etc. Although the computational cost may increase, this could pave the way to building at-scale simulations of mosquito infestations in the built environment. Weather data may be used to model the life cycle of each mosquito (see [51] for Cx. quinquefasciatus, [6] for Cx. pipiens and Cx. tarsalis, and [38] for Ae. aegypti); the tools of Appendix B can be used to digitize local maps showing buildings, trees, gardens, etc; and the forward model provides a way to simulate adult mosquitoes moving in this environment, with females being attracted to suitable oviposition sites when gravid.
@@ Acknowledgements
This material is partly based upon High Performance Computing (HPC) resources supported by the University of Arizona TRIF, UITS, and Research, Innovation, and Impact (RII) and maintained by the UArizona Research Technologies department. Research reported in this publication was supported in part by the University of Arizona’s BIO5 Institute Team Scholars Program. Mount Holyoke College undergraduate students, Chuhan Wang ‘24 and Karry Wang ‘23, worked on an image processing project in Summer 2022, which helped with the digitization of maps used in this work. The authors would like to thank Heidi Brown for helpful conversations on mosquito behavior. The authors would also like to thank the referees of the manuscript for their careful reading and helpful comments.
@@ Appendix A Diffusion Equation
As described in Section 2, diffusive mosquito dispersal means that at the microscopic level, the two-dimensional (ground) projection of the flight path of a moving mosquito consists of an alternation of straight flight segments of fixed length and angular reorientations uniformly distributed in $0,2\pi]$. At the macroscopic level, this corresponds to planar diffusion, described by the partial differential equation
@@ B.2 Processing: Release and recapture of diffusive agents
[…]
@@ C.1 Description of the method
and $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are 4-dimensional vectors of independent random numbers drawn from the uniform distribution in the interval $0,1]^{4}$. Note that “*” here denotes element-wise multiplication. The parameters are $w$, the “inertia” weight, $\phi_{p}$, the “cognitive” coefficient, and $\phi_{g}$, the “social” coefficient. These heuristic coefficients are chosen so that $0<w<1$ and $\phi_{p},\phi_{g}\geq 1$, reflecting a balance between moving towards the “best seen” position and swarming for a more effective search. In particular, with $w<1$, the first term acts to slow down the particle when getting close to a minimum. The remaining two terms move it generally in the direction of its own local best position (second term) and the swarm’s global best position (third term), with an element of randomness included in that motion. We initiate $\mathbf{v}$ randomly in $1,1]^{4}$, and denote the resulting vector from the previous step by $\mathbf{v}_{old}$. For all of the simulations presented here, we set $w=0.4$ and $\phi_{p}=\phi_{g}=1.2$. A new generation of particles is obtained after all of the positions of the previous generation have been updated once according to the above formula.
C.2 Computational cost
The code is not yet optimized for usability, but we believe its current computational cost is reasonable. Table 6 reports on the time elapsed (using the MATLAB tic and toc commands) to obtain an estimate of k from several runs of the RDA-PSO method. We report the average of k over multiple runs (performed in parallel to save computational time), as well as an interval corresponding to one standard deviation about the mean value of the k estimates. Simulations were run for the Hainan - center field experiment (left panel of figure 1). As usual, the parameter tuples associated with an error that was one standard deviation above the mean error were considered outliers and removed. Different configurations (laptop or high-performance computer (HPC), 2- or 4-parameter optimizations) were analyzed. Since the optimal value of p obtained for this study is close to zero, the 2-parameter and 4-parameter runs are expected to produce estimates of k that are close to one another, even though this is not generally the case.
Table 6: Elapsed time for 4-parameter (k, q, p, ${\tt s}_{\tt e}$) and 2-parameter (k, q) optimizations (with p=0 and ${\tt s}_{\tt e}$=1 for the latter), on a laptop with 8 workers and on an HPC using 1 node with 32 workers. The mean value of k and its range within one standard deviation are also reported.
number of elapsed time $\overline{{\tt k}}$ $\overline{{\tt k}}-\sigma,\>\overline{{\tt k}}+\sigma]$
RDA-PSO runs (in minutes)
4-parameter opt. laptop 1 10.53 36.99
10 38.65 36.93 [36.3, 37.56]
HPC 10 20.67 37.43 [36.34, 38.51]
100 69.93 36.95 [36.02, 37.89]
500 277.95 36.78 [35.83, 37.73]
2-parameter opt. laptop 10 33.17 36.46 [35.62, 37.29]
HPC 100 44.67 36.27 [35.58, 36.97]
[…]
For all of the configurations considered, the average of k is close to the one obtained using 500 RDA-PSO runs, which is the most accurate and requires the most computational time. One RDA-PSO run with 36 particles and 12 generations corresponds to 36*13 simulations, each of which is the average of 5 independent iterations of the forward model. So one RDA-PSO run corresponds to 2,340 release-recapture experiments. Looking at the elapsed time for 1 RDA-PSO run on the laptop (without parallelization) in Table 6, this is equivalent to 0.27 seconds per release-recapture experiment. The computational cost of each RDA-PSO run goes down as the number runs, executed in parallel, increases. For example, for 10 runs on the HPC, each run takes about 124 seconds. For 500 runs, each one takes about 33.35 seconds.
@@ Appendix D Virtual MRR experiment
This section illustrates how the different methods introduced in this article perform on a virtual MRR experiment. To this end, we first obtain synthetic capture ratios by simulating the forward problem with fixed parameters and averaging the resulting ratios of 100 simulations. This leads to very robust ratios that fully capture the observed dispersal. Consequently, all of the models, except for TC which uses uncorrected ratios, estimate values of $D$ that are reasonably close (of the order of 10% or less) to the known diffusion coefficient, thereby providing a validation of each of these approaches. The RDA-PSO method, iterated 100 times in this case, is highly accurate and provides the best results. Appendix G compares the different methods in the case of synthetic MRR experiments. By synthetic, we mean experiments that use realistic trap placements and observed ratios (obtained without averaging the forward model 100 times), but for which the diffusion coefficient is still known.
@@ D.2 Generation of MRR data for the virtual experiment
In order to limit the effect of randomness in the generated data, we run 100 simulations of the forward problem with the above setup and use average recapture ratios as the observed ratios. In each of the 100 simulations, N=N=100,000 agents are released from the middle of the grid. Each moves independently of the others and performs a random walk with parameters k = 100 and q = 30 for a total duration of 20 days. Trap efficiency is set at 100% and p = 0. Table 7 lists the average capture ratios, by day (with a count every other day) and by zone, resulting from these simulations. The selected value of k corresponds to a diffusion coefficient $D=300$ m2 per day.
@@
D.3 Application of the MDT model
We calculate the empirical MDT by multiplying the spatial ratios $\sigma_{j}$ element-wise by the average distance to each corresponding zone, that is